# Multiple ergodic averages along functions from a Hardy field: convergence, recurrence and combinatorial applications

### v2

@@ 4.   Characteristic Factors
In this section we show that nilsystems are characteristic for the ergodic averages (1.5). We will make use of the uniformity seminorms introduced in [HK05a] for ergodic systems. As was observed in [CFH11], the ergodicity of the system is not necessary.
@@
$\lvert\!|\!|h|\!|\!\rvert_{0}=\int_{X}h~{}\mathrm{d}\mu\qquad\text{and}\qquad\lvert\!|\!|h|\!|\!\rvert_{s}^{2^{s}}=\lim_{N\to\infty}\frac{1}{N}\sum_{n=1}^{N}\,\lvert\!|\!|\overline{h}\cdot T^{n}h|\!|\!\rvert_{s-1}^{2^{s-1}}\text{ forevery }s\in\mathbb{N}.$
The existence of the limits in this definition was established in [HK05a] for ergodic systems and in [CFH11, Section 2.2] in general.
@@
Let $\mathcal{H}$ be a Hardy field and $W\in\mathcal{H}$ with $1\prec W(t)\ll t$. Assume $f_{1},\ldots,f_{k}\in\mathcal{H}$ satisfy 1, $|f_{1}(t)|\ll\ldots\ll|f_{k}(t)|$, and $\lim_{t\to\infty}|f_{k}(t)|=\lim_{t\to\infty}|f_{k}(t)-f_{i}(t)|=\infty$ for every $i<k$. Then there exists $s\in\mathbb{N}$ such that for any invertible measure preserving system $(X,\mathcal{B},\mu,T)$ and any $h_{k}\in L^{\infty}(X)$ with $\lvert\!|\!|h_{k}|\!|\!\rvert_{s}=0$ we have
@@
4.1.   The sub-linear case of 4.2
For the proof of 4.2 we have to distinguish between the case when $f_{k}$ has at most linear growth (i.e. $|f_{k}(t)|\ll t$), and the case when $f_{k}$ has super-linear growth (i.e. $t\prec|f_{k}(t)|$), because the respective inductive procedures used to prove these two cases rely on different arguments. In this subsection we focus on the proof of the former case, which for the convenience of the reader we state as a separate theorem here. In this theorem we assume only that the “weight-function” $W(t)$ has sub-exponential growth (i.e., $W(t)\prec c^{t}$ for all $c>1$), in contrast to 4.2 where $W(t)$ is assumed to have at most linear growth (i.e., $W(t)\ll t$). This is because averages with sub-exponential weights show up naturally in our inductive procedure.
@@
Let $k\in\mathbb{N}$. Let $\mathcal{H}$ be a Hardy field and $W\in\mathcal{H}$ with $1\prec W(t)$ and suppose $W(t)$ has sub-exponential growth. Assume $f_{1},\ldots,f_{k}\in\mathcal{H}$ satisfy property 1, $|f_{1}(t)|\ll\ldots\ll|f_{k}(t)|\ll t$, and $\lim_{t\to\infty}|f_{k}(t)|=\lim_{t\to\infty}|f_{k}(t)-f_{i}(t)|=\infty$ for every $i<k$. Then there exists a constant $C_{k}>0$, depending only on $k$ and $f_{1},\dots,f_{k}$, such that for any invertible measure preserving system $(X,\mathcal{B},\mu,T)$ and any $h_{k}\in L^{\infty}(X)$ we have
$\sup_{h_{1},\ldots,h_{k-1}\in L^{\infty}}~{}\sup_{a\in\ell^{\infty}}~{}\left\|\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a(n)\prod_{i=1}^{k}T^{[f_{i}(n)]}h_{i}\right\|_{L^{2}}\leqslant\,C_{k}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}+{\rm o}_{N\to\infty}(1),$ (4.2)
@@
Given $f\colon\mathbb{R}\to\mathbb{R}$ and $g\colon\mathbb{R}\to(0,\infty)$, we write $f(t)\sim g(t)$ if $\lim_{t\to\infty}f(t)/g(t)=1$. The first step in proving 4.3 is to reduce it to the case $f_{k}(t)\sim ct$ for some $c>0$ using a method often described as “change of variables”.
@@
Suppose 4.3 has already been proven for a specific $k\in\mathbb{N}$ when $f_{k}(t)\sim ct$ for some $c>0$. Then 4.3 follows in full generality for this specific $k$.
@@
Proof of 4.4.
By replacing $f_{k}$ with $-f_{k}$ and $T$ with its inverse $T^{-1}$ if necessary, we can assume without loss of generality that $f_{k}(t)$ is eventually positive. If $f_{k}(t)\sim ct$ for some $c>0$ then there is nothing to show. Thus, we can assume $f_{k}(t)\prec t$. Let $g(t)\coloneqq f_{k}(t)$ and define $K_{j}\coloneqq\left\{n\in\mathbb{N}:j<g(n)\leqslant j+1\right\}.$ A straightforward calculation shows that for all but finitely many $j\in\mathbb{N}$ one has
@@
instead of (4.2) is suffices to show
$\sup_{h_{1},\ldots,h_{k-1}\in L^{\infty}}~{}\sup_{a\in\ell^{\infty}}~{}\left\|\frac{1}{W(N)}\sum_{j=1}^{\lfloor g(N)\rfloor}\sum_{n\in K_{j}}w(n)a(n)\prod_{i=1}^{k}T^{[f_{i}(n)]}h_{i}\right\|_{L^{2}}\leqslant\,C_{k}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}+{\rm o}_{N\to\infty}(1).$ (4.4)
Set $V=W\circ g^{-1}$. According to 4.5, $V(t)$ has sub-exponential growth. In particular, $\lim_{N\to\infty}\sup_{c\in[0,1]}\frac{V(N+c)}{V(N)}=1$, which implies that
@@
$\sup_{h_{1},\ldots,h_{k-1}\in L^{\infty}}~{}\sup_{a\in\ell^{\infty}}~{}\left\|\frac{1}{V(\lfloor g(N)\rfloor)}\sum_{j=1}^{\lfloor g(N)\rfloor}\sum_{n\in K_{j}}w(n)a(n)\prod_{i=1}^{k}T^{[f_{i}(n)]}h_{i}\right\|_{L^{2}}\leqslant\,C_{k}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}+{\rm o}_{N\to\infty}(1)$
@@
$\sup_{h_{1},\ldots,h_{k-1}\in L^{\infty}}~{}\sup_{a\in\ell^{\infty}}~{}\left\|\frac{1}{V(N)}\sum_{j=1}^{N}\sum_{n\in K_{j}}w(n)a(n)\prod_{i=1}^{k}T^{[f_{i}(n)]}h_{i}\right\|_{L^{2}}\leqslant\,C_{k}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}+{\rm o}_{N\to\infty}(1)$ (4.5)
Define $g_{i}(t)\coloneqq f_{i}(g^{-1}(t))$ for $i=1,\ldots,k$ and note that $g_{k}(t)=t$. Also note that $g_{1},\ldots,g_{k},V$ belong to the same Hardy field. Indeed, from $f_{1},\ldots,f_{k},W\in\mathcal{H}$ we conclude that $g_{1},\ldots,g_{k},V\in\mathcal{H}\circ g^{-1}$, and according to [Bos81, Lemma 6.4] if $g\in\mathcal{H}$ with $\lim_{t\to\infty}g(t)=\infty$ then $\mathcal{H}\circ g^{-1}=\{f\circ g^{-1}:f\in\mathcal{H}\}$ is a Hardy field. Since $|f_{i}(t)|\ll f_{k}(t)$ for all $i\leqslant k$, there exists $C\in\mathbb{N}$ such that $|f_{i}(t)|<Cf_{k}(t)$ for all $t$ in some half-line $[t_{0},\infty)$. This implies that $|f_{i}(n)-g_{i}(j)|<C$ for all $n\in K_{j}$ and all but finitely many $j\in\mathbb{N}$. In other words, for all but finitely many $j\in\mathbb{N}$ and all $n\in K_{j}$ we have $[f_{i}(n)]\in[g_{i}(j)]+\{-C,\ldots,C\}$. For every $\eta=(\eta_{1},\ldots,\eta_{k})\in\{-C,\ldots,C\}^{k}$ define
@@
$\sup_{h_{1},\ldots,h_{k-1}\in L^{\infty}}\,\sup_{a\in\ell^{\infty}}\left\|\frac{1}{V(N)}\sum_{j=1}^{N}\sum_{n\in K_{j}^{(\eta)}}w(n)a(n)\prod_{i=1}^{k}T^{[g_{i}(j)]}T^{\eta_{i}}h_{i}\right\|_{L^{2}}\leqslant\,C_{k,\eta}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}+{\rm o}_{N\to\infty}(1),$ (4.6)
for some constants $C_{k,\eta}$. Note that $\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}=\lvert\!|\!|T^{\eta_{k}}h_{k}|\!|\!\rvert_{k+1}$. Moreover, the supremum over all functions $h_{1},\ldots,h_{k-1}\in L^{\infty}(X)$ with $\|h_{i}\|_{L^{\infty}}\leqslant 1$ is the same as the supremum over all functions $T^{\eta_{1}}h_{1},\ldots,T^{\eta_{k-1}}h_{k-1}\in L^{\infty}(X)$ with $\|T^{\eta_{i}}h_{i}\|_{L^{\infty}}\leqslant 1$. This means that (4.6) holds if and only if for all $h_{k}\in L^{\infty}(X)$ we have
$\sup_{h_{1},\ldots,h_{k-1}\in L^{\infty}}~{}\sup_{a\in\ell^{\infty}}~{}\left\|\frac{1}{V(N)}\sum_{j=1}^{N}\sum_{n\in K_{j}^{(\eta)}}w(n)a(n)\prod_{i=1}^{k}T^{[g_{i}(j)]}h_{i}\right\|_{L^{2}}\leqslant\,C_{k,\eta}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}+{\rm o}_{N\to\infty}(1),$ (4.7)
@@
$\sup_{h_{1},\ldots,h_{k-1}\in L^{\infty}}~{}\sup_{b\in\ell^{\infty}}~{}\left\|\frac{1}{V(N)}\sum_{j=1}^{N}b(j)\left(\sum_{n\in K_{j}}w(n)\right)\prod_{i=1}^{k}T^{[g_{i}(j)]}h_{i}\right\|_{L^{2}}\leqslant\,C_{k,\eta}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}+{\rm o}_{N\to\infty}(1).$ (4.8)
@@
$\sup_{h_{1},\ldots,h_{k-1}\in L^{\infty}}~{}\sup_{b\in\ell^{\infty}}~{}\left\|\frac{1}{V(N)}\sum_{j=1}^{N}b(j)v(j)\prod_{i=1}^{k}T^{[g_{i}(j)]}h_{i}\right\|_{L^{2}}\leqslant\,C_{k,\eta}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}+{\rm o}_{N\to\infty}(1).$ (4.11)
Since $g_{k}(t)=t$, (4.11) follows from the hypothesis of the lemma once we have verified that $g_{1},\ldots,g_{k}$ satisfy property 1 with $W(t)$ replaced by $V(t)$. In other words, it remains to show that for all $f\in\nabla\text{-}\mathrm{span}(g_{1},\ldots,g_{k})$ and $p\in\mathbb{R}[t]$ either $|f(t)-p(t)|\ll 1$ or $|f(t)-p(t)|\succ\log(V(t))$. Since $g_{1},\ldots,g_{k}$ have at most linear growth, all it’s derivatives are either asymptotically constant, or asymptotically negligible. Therefore, it suffices to show the property for $f\in\mathrm{span}(g_{1},\ldots,g_{k})$ instead of $f\in\nabla\text{-}\mathrm{span}(g_{1},\ldots,g_{k})$. This follows readily from the fact that $f_{1},\ldots,f_{k}$ satisfy property 1 and $g_{i}=f_{i}\circ g^{-1}$ and $V=W\circ g^{-1}$. ∎
For the inductive step in the proof of 4.3 we need the following version of van der Corput’s Lemma.
@@
[…]
@@
$\begin{split}\lim_{k\to\infty}&\Bigg{\|}\frac{1}{H}\sum_{m=1}^{H}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}u_{N_{k}}(n+m)\Bigg{\|}^{2}\\ &\leqslant\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\left\|\frac{1}{H}\sum_{m=1}^{H}u_{N_{k}}(n+m)\right\|^{2}\\ &=\frac{1}{H^{2}}\sum_{m_{1},m_{2}=1}^{H}\left(\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\big{\langle}u_{N_{k}}(n+m_{1}),u_{N_{k}}(n+m_{2})\big{\rangle}\right)\\ &=\frac{1}{H^{2}}\sum_{1\leqslant m_{2}<m_{1}\leqslant H}2~{}{\rm Re}\left(\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\big{\langle}u_{N_{k}}(n+m_{1}),u_{N_{k}}(n+m_{2})\big{\rangle}\right)\,+\,{\rm O}\left(\frac{1}{H}\right)\\ &=\frac{1}{H^{2}}\sum_{1\leqslant m_{2}<m_{1}\leqslant H}2~{}{\rm Re}\left(\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\big{\langle}u_{N_{k}}(n+m_{1}-m_{2}),u_{N_{k}}(n)\big{\rangle}\right)\,+\,{\rm O}\left(\frac{1}{H}\right)\\ &={\rm Re}\left(\frac{1}{H}\sum_{m=1}^{H}\tfrac{2(H-m)}{H}\left(\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\big{\langle}u_{N_{k}}(n+m),u_{N_{k}}(n)\big{\rangle}\right)\right)\,+\,{\rm O}\left(\frac{1}{H}\right)\\ &\leqslant\left|\frac{1}{H}\sum_{m=1}^{H}\tfrac{2(H-m)}{H}\left(\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\big{\langle}u_{N_{k}}(n+m),u_{N_{k}}(n)\big{\rangle}\right)\right|\,+\,{\rm O}\left(\frac{1}{H}\right)\end{split}$
@@
$\sup_{a\in\ell^{\infty}}~{}\left\|\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a(n)T^{[f_{1}(n)]}h_{1}\right\|_{L^{2}}\leqslant\,C_{1}\lvert\!|\!|h_{1}|\!|\!\rvert_{2}+{\rm o}_{N\to\infty}(1),$ (4.15)
where the supremum is taken over all $a\in\ell^{\infty}(\mathbb{N})$ with $\|a\|_{\ell^{\infty}}\leqslant 1$. In light of 4.4, we can assume without loss of generality that $f_{1}(t)\sim ct$ for some $c>0$. Instead of taking the supremum in (4.15), it suffices to show that for any $a_{1},a_{2},\ldots\in\ell^{\infty}(\mathbb{N})$ with $\|a_{N}\|_{\ell^{\infty}}\leqslant 1$ we have
$\left\|\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a_{N}(n)T^{[f_{1}(n)]}h_{1}\right\|_{L^{2}}\leqslant\,C_{1}\lvert\!|\!|h_{1}|\!|\!\rvert_{2}+{\rm o}_{N\to\infty}(1).$ (4.16)
Note that $\lim_{N\to\infty}w(N)/W(N)=0$ because $W(t)$ has sub-exponential growth. By applying 4.6 with $W(N)=P_{N}$, $w(n)=p_{n}$, and $u_{N}(n)=a_{N}(n)T^{n}h_{1}$, we see that (4.16) follows if we can show
@@
Since $f_{1}(t)\sim ct$, we have $[f_{1}(n+m)]-[f_{1}(n)]\in[cm]+\{-2,-1,0,1,2\}$ for all but finitely many $n\in\mathbb{N}$. Define, for $\eta\in\{-2,-1,0,1,2\}$, the set $V_{\eta}\coloneqq\{n\in\mathbb{N}:[f_{1}(n+m)]-[f_{1}(n)]=[cm]+\eta\}$. Then, since $\bigcup_{\eta\in\{-2,-1,0,1,2\}}V_{\eta}$ is co-finite in $\mathbb{N}$, in place of (4.17) it suffices to show that for all $\eta\in\{-2,-1,0,1,2\}$,
@@
$\begin{split}\limsup_{H\to\infty}\,\limsup_{N\to\infty}~{}\frac{1}{H}\sum_{m=1}^{H}\Bigg{|}\int T^{[cm]+\eta}h_{1}\cdot&\overline{h}_{1}~{}\mathrm{d}\mu\Bigg{|}\\ &\left|\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a_{N}(n+m)\overline{a}_{N}(n)1_{V_{\eta}}(n)\right|\leqslant\,\frac{C_{1}}{5}\lvert\!|\!|h_{1}|\!|\!\rvert_{2}^{2},\end{split}$
@@
$\lim_{H\to\infty}~{}\frac{1}{H}\sum_{m=1}^{H}\left|\int T^{[cm]+\eta}h_{1}\cdot\overline{h}_{1}~{}\mathrm{d}\mu\right|\leqslant\,\frac{C_{1}}{5}\lvert\!|\!|h_{1}|\!|\!\rvert_{2}^{2}.$ (4.19)
@@
$\lim_{H\to\infty}\frac{1}{H}\sum_{m=1}^{H}\left|\int T^{m}h_{1}\cdot\overline{h}_{1}~{}\mathrm{d}\mu\right|~{}\leqslant~{}\lvert\!|\!|h_{1}|\!|\!\rvert_{2}^{2},$
by the definition of |||.$\lvert\!|\!|.|\!|\!\rvert_{2}$, and so (4.19) holds with $C_{1}=5(c+1)$.
For the proof of the inductive step, assume 4.3 has already been proven for $k-1$; we want to show
$\sup_{h_{1},\ldots,h_{k-1}}~{}\sup_{a\in\ell^{\infty}}~{}\left\|\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a(n)\prod_{i=1}^{k}T^{[f_{i}(n)]}h_{i}\right\|_{L^{2}}\leqslant\,C_{k}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}+{\rm o}_{N\to\infty}(1).$ (4.20)
Again, in light of 4.4 we can assume $f_{k}(t)\sim ct$ for some $c>0$, and instead of having the suprema $\sup_{h_{1},\ldots,h_{k-1}}$ and $\sup_{a\in\ell^{\infty}}$ in (4.20), it suffices to show that for any $h_{1,N},\ldots,h_{k-1,N}\in L^{\infty}(X)$ with $\|h_{i,N}\|_{L^{\infty}}\leqslant 1$ and any $a_{N}\in\ell^{\infty}(\mathbb{N})$ with $\|a_{N}\|_{\ell^{\infty}}\leqslant 1$ we have
$\left\|\frac{1}{W(N)}\sum_{n=1}^{N}w(n)a_{N}(n)\prod_{i=1}^{k}T^{[f_{i}(n)]}h_{i,N}\right\|_{L^{2}}\leqslant\,C_{k}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}+{\rm o}_{N\to\infty}(1),$ (4.21)
where, for convenience, we took $h_{k,N}=h_{k}$ for all $N\in\mathbb{N}$. We apply 4.6 once more, this time with $u_{N}(n)=a_{N}(n)\prod_{i=1}^{k}T^{[f_{i}(n)]}h_{i,N}$, and deduce that (4.21) holds if
@@
$\limsup_{H\to\infty}\limsup_{N\to\infty}\frac{1}{H}\sum_{m=1}^{H}\left\|\frac{1}{W(N)}\sum_{n=1}^{N}w(n)\tilde{a}_{m,N}(n)\prod_{i=1}^{k-1}T^{[\tilde{f}_{i}(n)]}\tilde{h}_{i,m,N}\right\|_{L^{2}}\leqslant\,\frac{C_{k}}{5^{k}3^{k}}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}^{2}.$ (4.25)
@@
$\limsup_{N\to\infty}\left\|\frac{1}{W(N)}\sum_{n=1}^{N}w(n)\tilde{a}_{m,N}(n)\prod_{i=1}^{k-1}T^{[\tilde{f}_{i}(n)]}\tilde{h}_{i,m,N}\right\|_{L^{2}}\leqslant\,C_{k-1}\lvert\!|\!|T^{[c_{k}m]+\eta_{k}}h_{k}\cdot\overline{h}_{k}|\!|\!\rvert_{k}.$
@@
$\limsup_{H\to\infty}\frac{1}{H}\sum_{m=1}^{H}C_{k-1}\lvert\!|\!|T^{[c_{k}m]+\eta_{k}}h_{k}\cdot\overline{h}_{k}|\!|\!\rvert_{k}\leqslant\,\frac{C_{k}}{5^{k}3^{k}}\lvert\!|\!|h_{k}|\!|\!\rvert_{k+1}^{2},$

### v3

@@ 4.  Characteristic Factors
In this section we show that nilsystems are characteristic for the ergodic averages (1.5). We will make use of the uniformity seminorms introduced in [HK05b] for ergodic systems. As was observed in [CFH11], the ergodicity of the system is not necessary.
@@
[…]
The existence of the limits in this definition was established in [HK05b] for ergodic systems and in [CFH11, Section 2.2] in general.
@@
[…]
@@
4.1.  The sub-linear case of ˜4.2
For the proof of ˜4.2 we have to distinguish between the case when $f_{k}$ has at most linear growth (i.e. $|f_{k}(t)|\ll t$), and the case when $f_{k}$ has super-linear growth (i.e. $t\prec|f_{k}(t)|$), because the respective inductive procedures used to prove these two cases rely on different arguments. In this subsection we focus on the proof of the former case, which for the convenience of the reader we state as a separate theorem here. In this theorem we assume only that the “weight-function” $W(t)$ has sub-exponential growth (i.e., $W(t)\prec c^{t}$ for all $c>1$), in contrast to ˜4.2 where $W(t)$ is assumed to have at most linear growth (i.e., $W(t)\ll t$). This is because averages with sub-exponential weights show up naturally in our inductive procedure.
@@
Let $k\in\mathbb{N}$. Let $\mathcal{H}$ be a Hardy field and $W\in\mathcal{H}$ with $1\prec W(t)$ and suppose $W(t)$ has sub-exponential growth. Assume $f_{1},\ldots,f_{k}\in\mathcal{H}$ satisfy property (P), $|f_{1}(t)|\ll\ldots\ll|f_{k}(t)|\ll t$, and $\lim_{t\to\infty}|f_{k}(t)|=\lim_{t\to\infty}|f_{k}(t)-f_{i}(t)|=\infty$ for every $i<k$. Then there exists a constant $C_{k}>0$, depending only on $k$ and $f_{1},\dots,f_{k}$, such that for any invertible measure preserving system $(X,\mathcal{B},\mu,T)$ and any $h_{k}\in L^{\infty}(X)$ we have
[…]
@@
Given $f\colon\mathbb{R}\to\mathbb{R}$ and $g\colon\mathbb{R}\to(0,\infty)$, we write $f(t)\sim g(t)$ if $\lim_{t\to\infty}f(t)/g(t)=1$. The first step in proving ˜4.3 is to reduce it to the case $f_{k}(t)\sim ct$ for some $c>0$ using a method often described as “change of variables”.
@@
Suppose ˜4.3 has already been proven for a specific $k\in\mathbb{N}$ when $f_{k}(t)\sim ct$ for some $c>0$. Then ˜4.3 follows in full generality for this specific $k$.
@@
Proof of ˜4.4.
By replacing $f_{k}$ with $f_{k}$ and $T$ with its inverse $T^{-1}$ if necessary, we can assume without loss of generality that $f_{k}(t)$ is eventually positive. If $f_{k}(t)\sim ct$ for some $c>0$ then there is nothing to show. Thus, we can assume $f_{k}(t)\prec t$. Let $g(t)\coloneqq f_{k}(t)$ and define $K_{j}\coloneqq\left\{n\in\mathbb{N}:j<g(n)\leqslant j+1\right\}.$ A straightforward calculation shows that for all but finitely many $j\in\mathbb{N}$ one has
@@
instead of (4.2) it suffices to show
[…]
Set $V=W\circ g^{-1}$. According to ˜4.5, $V(t)$ has sub-exponential growth. In particular, $\lim_{N\to\infty}\sup_{c\in[0,1]}\frac{V(N+c)}{V(N)}=1$, which implies that
@@
[…]
@@
[…]
Define $g_{i}(t)\coloneqq f_{i}(g^{-1}(t))$ for $i=1,\ldots,k$ and note that $g_{k}(t)=t$. Also note that $g_{1},\ldots,g_{k},V$ belong to the same Hardy field. Indeed, from $f_{1},\ldots,f_{k},W\in\mathcal{H}$ we conclude that $g_{1},\ldots,g_{k},V\in\mathcal{H}\circ g^{-1}$, and according to [Bos81, Lemma 6.4] if $g\in\mathcal{H}$ with $\lim_{t\to\infty}g(t)=\infty$ then $\mathcal{H}\circ g^{-1}=\{f\circ g^{-1}:f\in\mathcal{H}\}$ is a Hardy field. Since $|f_{i}(t)|\ll f_{k}(t)$ for all $i\leqslant k$, there exists $C\in\mathbb{N}$ such that $|f_{i}(t)|<Cf_{k}(t)$ for all $t$ in some half-line $t_{0},\infty)$. This implies that $|f_{i}(n)-g_{i}(j)|<C$ for all $n\in K_{j}$ and all but finitely many $j\in\mathbb{N}$. In other words, for all but finitely many $j\in\mathbb{N}$ and all $n\in K_{j}$ we have $f_{i}(n)]\in[g_{i}(j)]+\{-C,\ldots,C\}$. For every $\eta=(\eta_{1},\ldots,\eta_{k})\in\{-C,\ldots,C\}^{k}$ define
@@
[…]
@@
[…]
@@
[…]
Since $g_{k}(t)=t$, (4.11) follows from the hypothesis of the lemma once we have verified that $g_{1},\ldots,g_{k}$ satisfy property (P) with $W(t)$ replaced by $V(t)$. In other words, it remains to show that for all $f\in\nabla\text{-}\mathrm{span}(g_{1},\ldots,g_{k})$ and $p\in\mathbb{R}[t]$ either $|f(t)-p(t)|\ll 1$ or $|f(t)-p(t)|\succ\log(V(t))$. Since $g_{1},\ldots,g_{k}$ have at most linear growth, all their derivatives are either asymptotically constant, or asymptotically negligible. Therefore, it suffices to show the property for $f\in\mathrm{span}(g_{1},\ldots,g_{k})$ instead of $f\in\nabla\text{-}\mathrm{span}(g_{1},\ldots,g_{k})$. This follows readily from the fact that $f_{1},\ldots,f_{k}$ satisfy property (P) and $g_{i}=f_{i}\circ g^{-1}$ and $V=W\circ g^{-1}$. ∎
For the inductive step in the proof of ˜4.3 we need the following version of van der Corput’s Lemma.
@@
$\lim_{N\to\infty}P_{N}\,=\,\infty\qquad\text{and}\qquad\lim_{N\to\infty}\frac{p_{N}}{P_{N}}\,=\,0.$
@@
$\begin{split}\lim_{k\to\infty}&\Bigg\|\frac{1}{H}\sum_{m=1}^{H}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}u_{N_{k}}(n+m)\Bigg\|^{2}\\ &\leqslant\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\left\|\frac{1}{H}\sum_{m=1}^{H}u_{N_{k}}(n+m)\right\|^{2}\\ &=\frac{1}{H^{2}}\sum_{m_{1},m_{2}=1}^{H}\left(\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\big\langle u_{N_{k}}(n+m_{1}),u_{N_{k}}(n+m_{2})\big\rangle\right)\\ &=\frac{1}{H^{2}}\hskip-3.00003pt\sum_{1\leqslant m_{2}<m_{1}\leqslant H}\hskip-3.00003pt2~{\rm Re}\left(\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\big\langle u_{N_{k}}(n+m_{1}),u_{N_{k}}(n+m_{2})\big\rangle\right)\,+\,{\rm O}\left(\frac{1}{H}\right)\\ &=\frac{1}{H^{2}}\hskip-3.00003pt\sum_{1\leqslant m_{2}<m_{1}\leqslant H}\hskip-3.00003pt2~{\rm Re}\left(\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\big\langle u_{N_{k}}(n+m_{1}-m_{2}),u_{N_{k}}(n)\big\rangle\right)\,+\,{\rm O}\left(\frac{1}{H}\right)\\ &={\rm Re}\left(\frac{1}{H}\sum_{m=1}^{H}\tfrac{2(H-m)}{H}\left(\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\big\langle u_{N_{k}}(n+m),u_{N_{k}}(n)\big\rangle\right)\right)\,+\,{\rm O}\left(\frac{1}{H}\right)\\ &\leqslant\left|\frac{1}{H}\sum_{m=1}^{H}\tfrac{2(H-m)}{H}\left(\lim_{k\to\infty}\frac{1}{P_{N_{k}}}\sum_{n=1}^{N_{k}}p_{n}\big\langle u_{N_{k}}(n+m),u_{N_{k}}(n)\big\rangle\right)\right|\,+\,{\rm O}\left(\frac{1}{H}\right)\end{split}$
@@
[…]
where the supremum is taken over all $a\in\ell^{\infty}(\mathbb{N})$ with $\|a\|_{\ell^{\infty}}\leqslant 1$. In light of ˜4.4, we can assume without loss of generality that $f_{1}(t)\sim ct$ for some $c>0$. Instead of taking the supremum in (4.15), it suffices to show that for any $a_{1},a_{2},\ldots\in\ell^{\infty}(\mathbb{N})$ with $\|a_{N}\|_{\ell^{\infty}}\leqslant 1$ we have
[…]
Note that $\lim_{N\to\infty}w(N)/W(N)=0$ because $W(t)$ has sub-exponential growth. By applying ˜4.6 with $W(N)=P_{N}$, $w(n)=p_{n}$, and $u_{N}(n)=a_{N}(n)T^{n}h_{1}$, we see that (4.16) follows if we can show
@@
Since $f_{1}(t)\sim ct$, we have $f_{1}(n+m)]-[f_{1}(n)]\in[cm]+\{-2,-1,0,1,2\}$ for all but finitely many $n\in\mathbb{N}$. Define, for $\eta\in\{-2,-1,0,1,2\}$, the set $V_{\eta}\coloneqq\{n\in\mathbb{N}:[f_{1}(n+m)]-[f_{1}(n)]=[cm]+\eta\}$. Then, since $\bigcup_{\eta\in\{-2,-1,0,1,2\}}V_{\eta}$ is co-finite in $\mathbb{N}$, in place of (4.17) it suffices to show that for all $\eta\in\{-2,-1,0,1,2\}$,
@@
[…]
@@
[…]
@@
[…]
by the definition of $\lvert\!|\!|.|\!|\!\rvert_{2}$, and so (4.19) holds with $C_{1}=5(c+1)$.
For the proof of the inductive step, assume ˜4.3 has already been proven for k1k-1; we want to show
[…]
Again, in light of ˜4.4 we can assume $f_{k}(t)\sim ct$ for some $c>0$, and instead of having the suprema $\sup_{h_{1},\ldots,h_{k-1}}$ and $\sup_{a\in\ell^{\infty}}$ in (4.20), it suffices to show that for any $h_{1,N},\ldots,h_{k-1,N}\in L^{\infty}(X)$ with $\|h_{i,N}\|_{L^{\infty}}\leqslant 1$ and any $a_{N}\in\ell^{\infty}(\mathbb{N})$ with $\|a_{N}\|_{\ell^{\infty}}\leqslant 1$ we have
[…]
where, for convenience, we took $h_{k,N}=h_{k}$ for all $N\in\mathbb{N}$. We apply ˜4.6 once more, this time with $u_{N}(n)=a_{N}(n)\prod_{i=1}^{k}T^{[f_{i}(n)]}h_{i,N}$, and deduce that (4.21) holds if
@@
[…]
@@
[…]
@@
[…]